Please help me solve this:
Elizabeth is seven years older than Chloe. Four years ago, Elizabeth was twice as old as Chloe was. How old are they now?

Answer by Cromlix(4381)

Hi there,
Elizabeth is seven years older than Chloe.
E = 7 + C......(1)
Four years ago, Elizabeth was twice as old as Chloe was.
E - 4 = 2(C - 4).....(2)
Replace 'E' in Eq(2) with '7 + C' from Eq(1)
7 + C - 4 = 2(C - 4)
7 + C - 4 = 2C - 8
Collect like terms.
C - 2C = - 8 + 4 - 7
- C = - 11
Multiply both sides by -1
C = 11
Elizabeth is 18 years old
Chloe is 11 years old.
Hope this helps :-)